# The N-prime graph and the Subgroup Isomorphism Problem

### v1

@@ 5. Metacyclic subgroups of solvable groups: ˜D and ˜E
Our aim is then to show that the number above coincides with the total number of elements $y$ of order $q^{k}$ in $H$ with $\left\langle y\right\rangle\cap M=1$. So, let $y$ be such an element: observe that $y$ acts as an automorphism of order $q^{k}$ on the $t$-dimensional ${\rm{GF}}(q)$-vector space $M$, hence, up to conjugacy in ${\rm{GL}}_{t}(q)$, the action of $y$ is represented by an upper unitriangular matrix $Y$ in the canonical Jordan form. It is easily seen that the dimension of ${\bf C}_{M}(y)$ over ${\rm{GF}}(q)$ coincides with the number of Jordan blocks of $Y$. Denote by $I_{d}$, $J_{d}$ the $d$-dimensional identity matrix and the Jordan block with eigenvalue $1$ respectively, and set $N=J_{d}-I_{d}$. Then for a positive integer $h$ we get
$(J_{d})^{h}=\sum_{i=0}^{d-1}{{h}\choose{i}}N^{i}$
@@
and, in particular, we see that $(J_{d})^{q^{k}}$ contains the summand $N^{q^{k}}$ if $d>q^{k}$. Since in our situation $(J_{d})^{q^{k}}$ must be the identity matrix $I_{d}$ for every Jordan block of $Y$, we conclude that $d$ is at most $q^{k}$ for every block of $Y$. But now, recalling that $t=\dim M$ equals $q^{k}\dim{{\bf C}_{M}(y)}$ and that $\dim{{\bf C}_{M}(y)}$ is the number of Jordan blocks of $Y$, we conclude that the dimension of each block is precisely $q^{k}$.
@@
$I_{q^{k}}+J_{q^{k}}+(J_{q^{k}})^{2}+\cdots+(J_{q^{k}})^{q^{k}-1}=\sum_{i=0}^{q^{k}-1}\left(\sum_{j=i}^{q^{k}-1}{j\choose i}\right)N^{i},$
@@
(this equality, not difficult to check, is sometimes referred to as the Christmas Stocking Theorem), it is then clear that $I_{t}+Y+Y^{2}+\cdots+Y^{q^{k}-1}$ is a block-diagonal matrix whose blocks are all equal to the $q^{k}$-dimensional matrix having all zero entries except the top-right entry, which is $1$. The kernel of the endomorphism associated with this matrix has then a dimension equal to $t-\dim{\bf C}_{M}(y)$, and so it contains $n^{q^{k}-1}$ elements.
Taking into account that the above calculation yields the same result if we replace $y^{-1}$ by any generator of $\left\langle y\right\rangle$, we conclude that the number of elements $z$ of order $q^{k}$ lying in the Sylow $q$-subgroup $M\left\langle y\right\rangle$ of $H$ and with $\left\langle z\right\rangle\cap M=1$ is $n^{q^{k}-1}q^{k-1}(q-1)$. Multiplying this result by the number of Sylow $q$-subgroups of $H$, we obtain a total of $n^{q^{k}-1}pq^{k-1}(q-1)$ elements, including our element $b$. The desired conclusion now follows because this number coincides with the number of elements of order $q^{k}$ in $H$ that normalize some Sylow $p$-subgroup of $H$, and the proof is complete.
We are ready to prove a slightly stronger form of ˜D.
@@
we can find an element $b\in G$ such that $\varepsilon_{b}(v)\neq 0$ and $bN$ is conjugate to $yN$ in $G/N$. Note that, in particular, $|b|$ is a multiple of $q^{k}$, but in fact we have $|b|=q^{k}$ by [12, Theorem 2.3]. Now, let $u\in G$ be such that $bN=(yN)^{uN}$. Clearly $bN$ normalizes $\left\langle x^{u}N\right\rangle$, hence $x^{u},b\in G$ satisfy the following properties: $|x^{u}|=p$, $|b|=q^{k}$, the subgroup $H=\left\langle N,x^{u},b\right\rangle$ is such that $H/N\cong C_{p}\rtimes C_{q}^{k}$ is a Frobenius group, and $\varepsilon_{b}(v)\neq 0$.
If $r\neq q$, then, by the Schur-Zassenhaus Theorem, $N$ is complemented in $H$ by a Hall $\{p,q\}$-subgroup $H_{0}$, and we can choose $H_{0}$ so that $b\in H_{0}\cong H/N\cong T$. Now, if $a$ is an element of order $p$ in $H_{0}$, then the elements $a$ and $b$ satisfy the conclusions of the statement and $G$ is not a counterexample. On the other hand, if $r=q$, then we are in a position to apply Lemma˜5.1 to the subgroup $H$ (note that $\left\langle b\right\rangle\cap N=1$ because $|b|=|bN|=q^{k}$), and again $G$ is not a counterexample.
In view of the above discussion, we conclude that ${\bf F}(G)$ must be a $p$-group. By the proof of the main theorem in [11], we can choose an element $b\in G$ such that $|b|=q^{k}$ and $\varepsilon_{b}(v)\neq 0$; moreover, $b$ acts as an automorphism of order $q^{k}$ on ${\bf F}(G)$, because $G$ is solvable and therefore ${\bf C}_{G}({\bf F}(G))={\bf Z}({\bf F}(G))$. But then we may apply Lemma˜3.2, and again it turns out that $G$ is not a counterexample. This is the final contradiction that completes the proof.
@@
We first claim that $G^{\prime}$ is a $p$-group. Otherwise $G^{\prime}$ contains a subgroup $N$, normal in $G$, whose order is a prime $r\neq p$. Then, denoting by $\varphi:V\to V({\mathbb{Z}}(G/N))$ the natural homomorphism, we get $T\cap{\rm ker}(\varphi)=1$ and hence $\varphi(T)\cong T$. By the minimality of $G$, there are $a\in G^{\prime}$ and $g\in G$ such that $\left\langle\varphi(a),\varphi(g)\right\rangle=\left\langle\varphi(a)\right\rangle\rtimes\left\langle\varphi(g)\right\rangle\cong T$ and $\varphi(v)$ is conjugate to $\varphi(g)=gN$ in ${\mathbb{Q}}(G/N)$. On the other hand, by the main theorem of [5], the First Zassenhaus Conjecture holds for $G$; therefore $v$ is conjugate in ${\mathbb{Q}}G$ to an element $b$ of $G$, so that $bN$ and $gN$ are both conjugate to $\varphi(v)$ in ${\mathbb{Q}}(G/N)$. Since $gN$ and $bN$ are elements of $G/N$ that are conjugate in ${\mathbb{Q}}(G/N)$, as is well known, they are also conjugate in $G/N$, and hence we may assume $g=b$. Let $H=\left\langle N,a,b\right\rangle$. Then $H/N\cong T$. If $r\neq q$, then $G$ contains a Hall subgroup of the form $\left\langle a\right\rangle\rtimes\left\langle b\right\rangle\cong T$, a contradiction. Otherwise, since $N$ has order $q$ and $p$ is larger than $q$, we have $\left\langle N,a\right\rangle=N\times\left\langle a\right\rangle$. Recalling that $b$ is an element of order $q^{k}$, we have $\left\langle N,b\right\rangle=N\rtimes\left\langle b\right\rangle$ and so $H=(N\times\left\langle a\right\rangle)\rtimes\left\langle b\right\rangle$; now, as $\left\langle aN,bN\right\rangle$ is isomorphic to $T$, then so is $\left\langle a\right\rangle\rtimes\left\langle b\right\rangle$, against the assumption. This finishes the proof of our claim that $G^{\prime}$ is a $p$-group.
@@
Next, we consider two different situations, depending on whether ${\bf Z}(T)$ is trivial or not. Assume first ${\bf Z}(T)=1$, which is equivalent to the condition $u,v^{q^{k-1}}]\neq 1$. As above, we take $b$ in $G$ which is conjugate to $v$ in ${\mathbb{Q}}G$. If $u^{v}=u^{r}$, then $r\not\equiv 1\,\,{\rm mod}\,\,\,p$, therefore $u,v]=u^{r-1}$ and it has order $p^{h}$. By Remark˜5.3, the order of $G^{\prime}$ is a multiple of $p^{h}$ and therefore, by hypothesis, $b^{q^{k-1}}$ commutes with $G^{\prime}$, for otherwise there is $a\in G^{\prime}$ such that $\left\langle a\right\rangle\rtimes\left\langle g\right\rangle\cong T$ and $G$ would not be a counterexample. If $g\in G$, then $g^{b^{q^{k-1}}}=zg$ for some $z\in G^{\prime}$. Then $g=g^{b^{q^{k}}}=z^{q}g$ and therefore, $z=1$. This shows that $b^{q^{k-1}}$ is central in $G$ and hence so is $v^{q^{k-1}}$ in $V$, which contradicts the assumption $u,v^{q^{k-1}}]\neq 1$.
@@
Finally, assume that ${\bf Z}(T)\neq 1$. Since $T$ is nonabelian, we have that ${\bf Z}(T)$ has order $q^{t}$ for a suitable $0<t<k$; setting $z=v^{q^{k-t}}$, we have ${\bf Z}(T)=\left\langle z\right\rangle$. By the fact that the First Zassenhaus Conjecture holds for $G$, we can consider $x\in G$ and $\gamma\in{\mathbb{Q}}G$ such that $(uz)^{\gamma}=x$. Note also that the $q$-part $x_{q}$ of $x$ is $z^{\gamma}$, and the $p$-part $x_{p}$ is $u^{\gamma}$; since $u$ lies in $T^{\prime}\subseteq V^{\prime}$, in view of Remark˜5.3, we then have $x_{p}\in G^{\prime}$. Finally, let $b\in G$ and $\delta\in{\mathbb{Q}}G$ be such that $v^{\delta}=b$. Now, for any $y\in G$, we get
@@
(recall that the partial augmentation with respect to a given element $y\in G$ is a class function of ${\mathbb{Q}}G$). It immediately follows that $b^{q^{k-t}}$ is conjugate to $x_{q}$ in $G$, and we take $g\in G$ such that ${(b^{q^{k-t}})}^{g}=x_{q}$. Clearly the subgroup $H=\langle x_{p},b^{g}\rangle$ has a center of order at least $q^{t}$, but we claim that $|{\bf Z}(H)|$ is precisely $q^{t}$ (i.e., $H\cong T$). Set $Z={\bf O}_{q}({\bf Z}(H))$. We will first prove that $|Z|=q^{t}$, and then that ${\bf Z}(H)=Z$. Indeed, $Z$ centralizes a nontrivial subgroup of the cyclic $p$-group $G^{\prime}$ and hence it centralizes $G^{\prime}$; on the other hand, for $P\in{\rm Syl}_{p}(G)$, clearly $Z$ acts trivially on $P/G^{\prime}$ as well, thus $Z$ centralizes $P$. But as $Z$ lies in an (abelian) Hall $p^{\prime}$-subgroup of $G$, we conclude that $Z$ lies in ${\bf Z}(G)$. If now $|Z|$ is larger that $q^{t}$, then $b^{q^{k-t-1}}$ would lie in ${\bf Z}(G)$ and hence in ${\bf Z}({\mathbb{Q}}G)$. But then $v^{q^{k-t-1}}=(b^{q^{k-t-1}})^{\delta^{-1}}=b^{q^{k-t-1}}$ would be central in $V$, contradicting the fact that $v^{q^{k-t-1}}$ does not centralize $u$. Therefore $|Z|=q^{t}$. As $t<k$ and $b^{g}$ has order $q^{k}$, $H$ is not abelian and hence $Z(H)=Z$ and $H\cong T$.
Observe that, if the elements $u$ and $v$ in the above statement generate an abelian subgroup of order $p^{h}q^{k}$, then $G$ has a subgroup isomorphic to $\left\langle u,v\right\rangle$ by [11]. This, together with Theorem 5.4, completes the proof of ˜E as stated in the introduction.
Acknowledgment. This work has been done during a visit of the second author at Dipartimento di Matematica e Informatica (DIMAI) of Università degli Studi di Firenze, funded by grant MICIU/PRX23/00370. He wishes to thank DIMAI for the hospitality and the Spanish Ministerio de Ciencia, Innovación y Universidades for the financial support.

### v2

@@ 5. Metacyclic subgroups of solvable groups: Theorem D and Theorem E
Our aim is then to show that the number above coincides with the total number of elements $y$ of order $q^{k}$ in $H$ with $\left\langle y\right\rangle\cap M=1$. So, let $y$ be such an element: observe that $y$ acts as an automorphism of order $q^{k}$ on the ${\rm{GF}}(q)$-vector space $M$, whose dimension we denote by $t$. Hence, up to conjugacy in ${\rm{GL}}_{t}(q)$, the action of $y$ is represented by an upper unitriangular matrix $Y$ in the canonical Jordan form. It is easily seen that the dimension of ${\bf C}_{M}(y)$ over ${\rm{GF}}(q)$ coincides with the number of Jordan blocks of $Y$. Denote by $I_{d}$, $J_{d}$ the $d$-dimensional identity matrix and the Jordan block with eigenvalue $1$ respectively, and set $E=J_{d}-I_{d}$. Then for a positive integer $h$ we get
$(J_{d})^{h}=\sum_{i=0}^{d-1}{{h}\choose{i}}E^{i}$
@@
and, in particular, we see that $(J_{d})^{q^{k}}$ contains the summand $E^{q^{k}}$ if $d>q^{k}$. Since in our situation $(J_{d})^{q^{k}}$ must be the identity matrix $I_{d}$ for every Jordan block of $Y$, we conclude that $d$ is at most $q^{k}$ for every block of $Y$. But now, recalling that $t=\dim M$ equals $q^{k}\dim{{\bf C}_{M}(y)}$ and that $\dim{{\bf C}_{M}(y)}$ is the number of Jordan blocks of $Y$, we conclude that the dimension of each block is precisely $q^{k}$.
@@
$I_{q^{k}}+J_{q^{k}}+(J_{q^{k}})^{2}+\cdots+(J_{q^{k}})^{q^{k}-1}=\sum_{i=0}^{q^{k}-1}\left(\sum_{j=i}^{q^{k}-1}{j\choose i}\right)E^{i},$
@@
(this equality, not difficult to check, is sometimes referred to as the Christmas Stocking Theorem), it is then clear that $I_{t}+Y+Y^{2}+\cdots+Y^{q^{k}-1}$ is a block-diagonal matrix whose blocks are all equal to the $q^{k}$-dimensional matrix having all zero entries except the top-right entry, which is $1$. The kernel of the endomorphism associated with this matrix has then a dimension equal to $t-\dim{\bf C}_{M}(y)$, and so it contains $\ell^{q^{k}-1}$ elements.
Taking into account that the above calculation yields the same result if we replace $y^{-1}$ by any generator of $\left\langle y\right\rangle$, we conclude that the number of elements $z$ of order $q^{k}$ lying in the Sylow $q$-subgroup $M\left\langle y\right\rangle$ of $H$ and with $\left\langle z\right\rangle\cap M=1$ is $\ell^{q^{k}-1}q^{k-1}(q-1)$. Multiplying this result by the number of Sylow $q$-subgroups of $H$, we obtain a total of $\ell^{q^{k}-1}pq^{k-1}(q-1)$ elements, including our element $b$. The desired conclusion now follows because this number coincides with the number of elements of order $q^{k}$ in $H$ that normalize some Sylow $p$-subgroup of $H$, and the proof is complete.
We are ready to prove a slightly stronger form of Theorem D.
@@
we can find an element $b\in G$ such that $\varepsilon_{b}(v)\neq 0$ and $bN$ is conjugate to $yN$ in $G/N$. Note that, in particular, $|b|$ is a multiple of $q^{k}$, but in fact we have $|b|=q^{k}$ by [13, Theorem 2.3]. Now, let $g\in G$ be such that $bN=(yN)^{gN}$. Clearly $bN$ normalizes $\left\langle x^{g}N\right\rangle$, hence $x^{g},b\in G$ satisfy the following properties: $|x^{g}|=p$, $|b|=q^{k}$, the subgroup $H=\left\langle N,x^{g},b\right\rangle$ is such that $H/N\cong C_{p}\rtimes C_{q}^{k}$ is a Frobenius group, and $\varepsilon_{b}(v)\neq 0$.
If $r\neq q$, then, by the Schur-Zassenhaus Theorem, $N$ is complemented in $H$ by a Hall $\{p,q\}$-subgroup $H_{0}$, and we can choose $H_{0}$ so that $b\in H_{0}\cong H/N\cong T$. Now, if $a$ is an element of order $p$ in $H_{0}$, then the elements $a$ and $b$ satisfy the conclusions of the statement and $G$ is not a counterexample. On the other hand, if $r=q$, then we are in a position to apply Lemma 5.1 to the subgroup $H$ (note that $\left\langle b\right\rangle\cap N=1$ because $|b|=|bN|=q^{k}$), and again $G$ is not a counterexample.
In view of the above discussion, we conclude that ${\bf F}(G)$ must be a $p$-group. By the main theorem in [12], we can choose an element $b\in G$ such that $|b|=q^{k}$ and $\varepsilon_{b}(v)\neq 0$; moreover, $b$ acts as an automorphism of order $q^{k}$ on ${\bf F}(G)$, because $G$ is solvable and therefore ${\bf C}_{G}({\bf F}(G))={\bf Z}({\bf F}(G))$. But then we may apply Lemma 3.2, and again it turns out that $G$ is not a counterexample. This is the final contradiction that completes the proof.
@@
We first claim that $G^{\prime}$ is a $p$-group. Otherwise $G^{\prime}$ contains a subgroup $N$, normal in $G$, whose order is a prime $r\neq p$. Then, denoting by $\varphi:V\to V({\mathbb{Z}}(G/N))$ the natural homomorphism, we get $T\cap{\rm ker}(\varphi)=1$ and hence $\varphi(T)\cong T$. By the minimality of $G$, there are $a\in G^{\prime}$ and $g\in G$ such that $\left\langle\varphi(a),\varphi(g)\right\rangle=\left\langle\varphi(a)\right\rangle\rtimes\left\langle\varphi(g)\right\rangle\cong T$ and $\varphi(v)$ is conjugate to $\varphi(g)=gN$ in ${\mathbb{Q}}(G/N)$. On the other hand, by the main theorem of [5], the First Zassenhaus Conjecture holds for $G$; therefore $v$ is conjugate in ${\mathbb{Q}}G$ to an element $b$ of $G$, so that $bN$ and $gN$ are both conjugate to $\varphi(v)$ in ${\mathbb{Q}}(G/N)$. Since $gN$ and $bN$ are elements of $G/N$ that are conjugate in ${\mathbb{Q}}(G/N)$, every irreducible character of $G/N$ takes the same value on $bN$ and $gN$. Hence they are in fact conjugate in $G/N$, and we may assume $g=b$. Let $H=\left\langle N,a,b\right\rangle$. Then $H/N\cong T$. If $r\neq q$, then $G$ contains a Hall subgroup of the form $\left\langle a\right\rangle\rtimes\left\langle b\right\rangle\cong T$, a contradiction. Otherwise, since $N$ has order $q$ and $p$ is larger than $q$, we have $\left\langle N,a\right\rangle=N\times\left\langle a\right\rangle$. Recalling that $b$ is an element of order $q^{k}$, we have $\left\langle N,b\right\rangle=N\rtimes\left\langle b\right\rangle$ and so $H=(N\times\left\langle a\right\rangle)\rtimes\left\langle b\right\rangle$; now, as $\left\langle aN,bN\right\rangle$ is isomorphic to $T$, then so is $\left\langle a\right\rangle\rtimes\left\langle b\right\rangle$, against the assumption. This finishes the proof of our claim that $G^{\prime}$ is a $p$-group.
@@
Next, we consider two different situations, depending on whether ${\bf Z}(T)$ is trivial or not. Assume first ${\bf Z}(T)=1$, which is equivalent to the condition $u,v^{q^{k-1}}]\neq 1$. As above, we take $b$ in $G$ which is conjugate to $v$ in ${\mathbb{Q}}G$. If $u^{v}=u^{r}$, then $r\not\equiv 1\,\,{\rm mod}\,\,\,p$, therefore $u,v]=u^{r-1}$ and it has order $p^{h}$. By Remark 5.3, the order of $G^{\prime}$ is a multiple of $p^{h}$ and therefore, by hypothesis, $b^{q^{k-1}}$ commutes with $G^{\prime}$, for otherwise there is $a\in G^{\prime}$ such that $\left\langle a\right\rangle\rtimes\left\langle b\right\rangle\cong T$ and $G$ would not be a counterexample. If $g\in G$, then $g^{b^{q^{k-1}}}=zg$ for some $z\in G^{\prime}$. Then $g=g^{b^{q^{k}}}=z^{q}g$ and therefore, $z=1$. This shows that $b^{q^{k-1}}$ is central in $G$ and hence so is $v^{q^{k-1}}$ in $V$, which contradicts the assumption $u,v^{q^{k-1}}]\neq 1$.
@@
Finally, assume that ${\bf Z}(T)\neq 1$. Since $T$ is nonabelian, we have that ${\bf Z}(T)$ has order $q^{t}$ for a suitable $0<t<k$; setting $z=v^{q^{k-t}}$, we have ${\bf Z}(T)=\left\langle z\right\rangle$. By the fact that the First Zassenhaus Conjecture holds for $G$, we can consider $x\in G$ and $\gamma\in{\mathbb{Q}}G$ such that $(uz)^{\gamma}=x$. Note also that the $q$-part $x_{q}$ of $x$ is $z^{\gamma}$, and the $p$-part $x_{p}$ is $u^{\gamma}$; since $u$ lies in $T^{\prime}\subseteq V^{\prime}$, in view of Remark 5.3, we then have $x_{p}\in G^{\prime}$. Finally, let $b\in G$ and $\delta\in{\mathbb{Q}}G$ be such that $v^{\delta}=b$. Now, for any $y\in G$, we get
@@
(recall that the partial augmentation with respect to a given element $y\in G$ is a class function of ${\mathbb{Q}}G$). It immediately follows that $b^{q^{k-t}}$ is conjugate to $x_{q}$ in $G$, and we take $g\in G$ such that ${(b^{q^{k-t}})}^{g}=x_{q}$. Clearly the subgroup $H=\langle x_{p},b^{g}\rangle$ has a center of order at least $q^{t}$, but we claim that $|{\bf Z}(H)|$ is precisely $q^{t}$ (i.e., $H\cong T$). Set $Z={\bf O}_{q}({\bf Z}(H))$. We will first prove that $|Z|=q^{t}$, and then that ${\bf Z}(H)=Z$. Indeed, $Z$ centralizes a nontrivial subgroup of the cyclic $p$-group $G^{\prime}$ and hence it centralizes $G^{\prime}$; on the other hand, for $P\in{\rm Syl}_{p}(G)$, clearly $Z$ acts trivially on $P/G^{\prime}$ as well, thus $Z$ centralizes $P$. But as $Z$ lies in an (abelian) Hall $p^{\prime}$-subgroup of $G$, we conclude that $Z$ lies in ${\bf Z}(G)$. If now $|Z|$ is larger that $q^{t}$, then $b^{q^{k-t-1}}$ would lie in ${\bf Z}(G)$ and hence in ${\bf Z}({\mathbb{Q}}G)$. But then $v^{q^{k-t-1}}=(b^{q^{k-t-1}})^{\delta^{-1}}=b^{q^{k-t-1}}$ would be central in $V$, contradicting the fact that $v^{q^{k-t-1}}$ does not centralize $u$. Therefore $|Z|=q^{t}$. As $t<k$ and $b^{g}$ has order $q^{k}$, $H$ is not abelian and hence ${\bf Z}(H)=Z$ and $H\cong T$.
Observe that, if the elements $u$ and $v$ in the above statement generate an abelian subgroup of order $p^{h}q^{k}$, then $G$ has a subgroup isomorphic to $\left\langle u,v\right\rangle$ by [12]. This, together with Theorem 5.4, completes the proof of Theorem E as stated in the introduction.
Acknowledgment. This work has been done during a visit of the second author at Dipartimento di Matematica e Informatica (DIMAI) of Università degli Studi di Firenze, funded by grant MICIU/PRX23/00370. He wishes to thank DIMAI for the hospitality and the Spanish Ministerio de Ciencia, Innovación y Universidades for the financial support. The authors also wish to thank Leo Margolis for his careful reading and useful comments on this paper.